- Remote participants by eZuce: None
Report of work (*:responsible person)
- Alignment: *Enomoto
- PD in the BS chambers was moved to REFL table.
- Offset of coil was offloaded into motors upstream.
- AOS: *Michimura, Yokogawa, Sugimono
- Alignment of green light to ETMY was done, and beam was returned to center area.
- CRY: mirror temperature increase from 24K 25K during night.
Work plan (*:responsible person)
- AOS: *Michimura
- Alignment of green at the ETMY table will be done today, so beam can be used outside of ETMY chamber around the ETMY area. However no laser hazard is needed at the ETMY area because power is very small.
- Tcam: *Inoue, Cory, Binhua
- Illuminator or IYA will be implemented this afternoon.
- Maintenance of PCs will be done tomorrow. DAQ and some RT models will be stopped.
- [AM] IR: PSL and Y-end, Green: PR2.
- [PM] IR: upstream SR2-3 line and Y-end, Green: PR2
- week of 3/5
- 3/5 Exchanging items and cleaning inside the PR2 chamber
- 3/8 IYC door will be closed.
- week of 3/12
- 3/14: meeting for pump-down check
- TMS installation at ETMY.
- Connecting ducts at the corner area.
- week of 3/19
- Vacuum test at at the corner area.
- TMS alignment using green light at ETMY.
- Beam alignment to REFL port.
- Connection of PD signal to DAQ should be implemented this week or next week.
- GigE camera should be placed at the REFL port eventually.